### v1

# MVG-CRPS: A Robust Loss Function for Multivariate Probabilistic Forecasting

Abstract
In probabilistic time series forecasting, the multivariate Gaussian (MVG) distribution is widely used as predictive distribution for correlated continuous random variables. Current deep probabilistic models typically employ neural networks to parameterize the mean vector and covariance matrix of the distribution, with log-score (i.e., negative log-likelihood) as the default loss function. However, log-score is highly sensitive to outliers, leading to significant errors when anomalies are present in the data. Motivated by the use of the continuous ranked probability score (CRPS) in learning univariate distributions, we propose a robust loss function specifically designed for high-dimensional MVG outputs. The proposed MVG-CRPS loss function has a closed-form expression based on the neural network outputs, making it easily integrable into deep learning models. We evaluate MVG-CRPS on two probabilistic forecasting tasks—multivariate autoregressive and univariate sequence-to-sequence (Seq2Seq) forecasting—both involving observations following MVG distribution. Experimental results on real-world datasets demonstrate that MVG-CRPS achieves both robustness and efficiency, offering enhanced accuracy and uncertainty quantification in probabilistic forecasting.
1 INTRODUCTION
Probabilistic forecasting aims to capture the uncertainty inherent in time series data, offering a distribution of possible future values rather than single-point estimates provided by deterministic forecasts. This approach is invaluable in various domains such as finance (Groen et al.,, 2013), weather prediction (Palmer,, 2012), and health care management (Jones and Spiegelhalter,, 2012), where understanding the range of potential outcomes is crucial for risk assessment and informed decision-making.
To achieve accurate probabilistic forecasts, it is crucial to develop a well-specified forecasting model and employ an effective evaluation metric during training. While the development of probabilistic forecasting models has received significant attention in recent years, the advancement of evaluation metrics has been less explored (Bjerregård et al.,, 2021). A common approach for evaluating a forecaster involves applying scoring rules that assign a numerical score to each forecast, which are then collected across a test dataset. The choice of scoring rule typically depends on the type of forecasting task—whether it is deterministic or probabilistic, univariate or multivariate—and the nature of the variable being forecasted, such as categorical or continuous. In this paper, we focus on continuous variables within the probabilistic forecasting setting.
Although scoring rules for univariate probabilistic forecasting are well-established, their multivariate counterparts remain less developed (Panagiotelis et al.,, 2023). Multivariate time series data, involving multiple interdependent variables, present additional challenges such as increased complexity, higher dimensionality, and intricate cross-variable interactions. Traditional evaluation metrics like the CRPS (Matheson and Winkler,, 1976) and the log-score have been effective in univariate contexts but often fall short in multivariate scenarios due to issues like sensitivity to outliers and computational inefficiency. For example, when extended to the multivariate case, the CRPS lacks a closed-form expression, leading to computational inefficiency; additionally, the log-score heavily penalizes the tails of the data distribution, making it more sensitive to outliers (Gebetsberger et al.,, 2018; Bjerregård et al.,, 2021).
To address these challenges, we propose a robust loss function specifically designed for multivariate probabilistic time series forecasting with a Gaussian distribution output layer. Our method utilizes a PCA whitening transformation to decorrelate the multivariate time series variables into a new random vector with zero mean and identity covariance matrix. Consequently, each component of this vector follows a standard Gaussian distribution, allowing us to apply the closed-form expression of the CRPS for evaluation. This approach not only overcomes the lack of a closed-form CRPS in the multivariate case but also reduces sensitivity to outliers and extreme tails, enhancing both robustness and computational efficiency. The key contributions of our work are:
  • We introduce a loss function that is less sensitive to outliers and extreme tails of the data distribution. This robustness ensures that the forecaster focuses on the overall data structure rather than being unduly influenced by anomalies.
  • The proposed loss function has a closed-form expression, allowing for the analytical computation of derivatives. This feature is particularly advantageous for training deep learning models, as it enables seamless integration with backpropagation algorithms and reduces computational overhead.
  • We conduct extensive experiments using deep probabilistic forecasting models on real-world datasets. Our results show that the proposed loss function outperforms traditional scoring rules in terms of both accuracy and robustness.
2 RELATED WORK
2.1 Probabilistic Forecasting
Probabilistic forecasting concentrates on modeling the probability distribution of target variables, as opposed to deterministic forecasting, which offers only single-point estimates. This approach is vital for capturing the uncertainty inherent in time series data, enabling better risk assessment and decision-making. There are two primary methodologies for probabilistic forecasting: parametric (e.g., through probability density functions (PDFs)) and non-parametric (e.g., through quantile functions) (Benidis et al.,, 2022).
PDF-based methods involve assuming a specific probability distribution (e.g., Gaussian, Poisson) and using neural networks to estimate the parameters of that distribution. For example, DeepAR (Salinas et al.,, 2020) utilizes a recurrent neural network (RNN) to model hidden state transitions and predict distribution parameters at each time step. Its multivariate extension, GPVar (Salinas et al.,, 2019), employs a Gaussian copula to convert observations into Gaussian variables, thereby assuming a joint MVG distribution. This technique captures dependencies among multiple time series, which is crucial for accurate multivariate probabilistic forecasting. In the Gaussian case, one can further introduce a batch dimension and use Generalized Least Square (GLS) method to capture the autocorrelation and cross-correlation among different time steps (Zheng et al.,, 2024; Zheng and Sun,, 2024).
Neural networks can also generate parameters for more complex probabilistic models. The deep state space model (Rangapuram et al.,, 2018) employs an RNN to produce parameters for state space models, effectively modeling both temporal dynamics and forecast uncertainty. The normalizing Kalman filter (de Bézenac et al.,, 2020) integrates normalizing flows with linear Gaussian state space models to handle nonlinear dynamics and compute the PDFs of observations. In this framework, RNNs generate state space model parameters at each time step, and normalizing flows transform the output into observations, allowing for more flexible distribution modeling. Normalizing flows (Rasul et al., 2021b, ) model complex, non-Gaussian distributions by transforming a simple base distribution through a sequence of invertible functions. Diffusion models (Rasul et al., 2021a, ) offer another strategy by defining a stochastic process that incrementally adds noise to the data and learning to reverse this process to generate samples from the desired distribution. Copula-based methods have also been explored to model dependencies between multiple time series. Studies by Drouin et al., (2022) and Ashok et al., (2024) use copulas to construct multivariate distributions by specifying individual marginal distributions and a copula function that captures the dependence structure. This allows for flexible modeling of inter-variable relationships, which is essential in multivariate probabilistic forecasting.
Quantile-based approaches focus on predicting specific quantiles of the target distribution without relying on strict parametric assumptions. For instance, MQ-RNN (Wen et al.,, 2017) produces quantile forecasts using a Seq2Seq RNN architecture. By forecasting multiple quantiles, these models can reconstruct the entire target distribution, which is particularly useful for capturing asymmetric or heavy-tailed behaviors.
Among the existing works that model the full probability distribution using a parametric approach, most employ the log-score as the loss function for optimizing model parameters.
2.2 Scoring Rules
Scoring rules are fundamental tools used to evaluate the quality of probabilistic forecasts by assigning numerical scores based on the predicted probability distributions and the observed outcomes. They not only assess the accuracy and calibration of forecasts but can also be utilized in the estimation of model parameters during training. A scoring rule is considered proper if the expected score is minimized when the forecasted probability distribution $p$ matches the true distribution $q$ of the observations. Formally, a scoring rule $s(p,q)$ is proper if the divergence $d(p,q)=s(p,q)-s(q,q)$ is non-negative and it is strictly proper if $d(p,q)=0$ implies $p=q$ (Bröcker,, 2009).
One of the simplest and most widely used scoring rules in both univariate and multivariate settings is the log-score, also known as the negative log-likelihood (NLL). It is defined as the negative logarithm of the predictive density evaluated at the observed value. This score is particularly common when a parametric form of the predictive density is available (Panagiotelis et al.,, 2023). The log-score is a strictly proper scoring rule and has several desirable properties, such as consistency and sensitivity to the entire distribution. However, it is known to lack robustness because it heavily penalizes forecasts that assign low probability to the observed outcome, making it highly sensitive to outliers and extreme events (Gneiting et al.,, 2007).
To address the sensitivity of the log-score to outliers, CRPS has been proposed as a more robust alternative, especially in univariate settings. CRPS measures the difference between the cumulative distribution functions (CDFs) of the forecast and the observation, effectively integrating the absolute error over all possible threshold values. It can be interpreted as a generalized version of the mean absolute error (MAE) (Gneiting et al.,, 2005). A key distinction between the log-score and CRPS is that CRPS grows linearly with the normalized prediction error, whereas the log-score grows quadratically. Consequently, the log-score assigns harsher penalties to poor probabilistic forecasts, increasing its sensitivity to outliers.
CRPS has been effectively used for parameter estimation in statistical models. For example, the method of minimum CRPS estimation has been introduced for fitting ensemble model output statistics (EMOS) coefficients, optimizing the forecasts by directly minimizing the CRPS rather than maximizing the likelihood (Gneiting et al.,, 2005). This approach can lead to better-calibrated predictive distributions that are not excessively wide or overdispersed.
Extending scoring rules to multivariate probabilistic forecasting introduces additional complexity due to the dependencies among variables and higher dimensionality. While the log-score remains applicable in multivariate contexts when a parametric form of the joint predictive density is available, it continues to suffer from sensitivity to outliers.
The Energy Score (Gneiting and Raftery,, 2007) is a commonly used multivariate generalization of the CRPS. It assesses the discrepancy between the forecasted and observed distributions by considering the expectations of distances between random vectors. It is reported that the energy score is effective at detecting errors in the mean of the forecast distribution but is less sensitive to errors in the variance and, more critically, to misspecifications in the correlation structure among variables (Pinson and Tastu,, 2013; Alexander et al.,, 2024). The absence of a closed form expression also necessitates the use of Monte Carlo simulations to approximate the score by drawing samples from the predictive distribution, which can be computationally intensive (Panagiotelis et al.,, 2023).
Other multivariate scoring rules include the Variogram Score (Scheuerer and Hamill,, 2015) and the Dawid-Sebastiani Score (Wilks,, 2020). The variogram score focuses on assessing the spatial or temporal dependence structure by comparing differences between pairs of observations and forecasts. The Dawid-Sebastiani score combines aspects of both the log-score and variance to evaluate forecasts. However, like the energy score, these metrics can face computational challenges and may require approximations when closed-form solutions are unavailable. We refer readers to the comprehensive reviews by Gneiting and Katzfuss, (2014) and Tyralis and Papacharalampous, (2024) for further details.
3 OUR METHOD
3.1 Multivariate Probabilistic Forecasting
Probabilistic forecasting aims to estimate the joint distribution over a collection of future quantities based on a given history of observations (Gneiting and Katzfuss,, 2014). Denote the time series vector at a time point $t$ as $\mathbf{z}_{t}=\left[z_{1,t},\dots,z_{N,t}\right]^{\top}\in\mathbb{R}^{N}$, where $N$ is the number of series. The problem of probabilistic forecasting can be formulated as $p\left(\mathbf{z}_{{T+1}:{T+Q}}\mid\mathbf{z}_{{T-P+1}:{T}};\mathbf{x}_{{T-P+1}:{T+Q}}\right)$, where $\mathbf{z}_{{t_{1}}:{t_{2}}}=\left[\mathbf{z}_{t_{1}},\ldots,\mathbf{z}_{t_{2}}\right]$, $P$ is the conditioning range, $Q$ is the prediction range, and $T$ is the time point that splits the conditioning range and prediction range. $\mathbf{x}_{t}$ are some known covariates for both past and future time steps.
Multivariate probabilistic forecasting can be formulated in different ways. One way is over the time series dimension, where multiple interrelated variables are forecasted simultaneously at each time point. Considering an autoregressive model, where the predicted output is used as input for the next time step, this formulation can be factorized as
$p\left(\mathbf{z}_{{T+1}:{T+Q}}\mid\mathbf{z}_{{T-P+1}:{T}};\mathbf{x}_{{T-P+1}:{T+Q}}\right)$ (1)
$=$ $\prod_{t=T+1}^{T+Q}p\left(\mathbf{z}_{t}\mid\mathbf{z}_{{t-P}:{t-1}};\mathbf{x}_{{t-P}:{t}}\right)=\prod_{t=T+1}^{T+Q}p\left(\mathbf{z}_{t}\mid\mathbf{h}_{t}\right),$
where $\mathbf{h}_{t}$ is a state vector that encodes all the conditioning information used to generate the distribution parameters, typically via a neural network.
Another option is over the prediction horizon, where forecasts are made across multiple future time steps for one or more variables, capturing temporal dependencies and uncertainties over time. Considering a shared model across different series:
$p\left(\mathbf{z}_{i,{T+1}:{T+Q}}\mid\mathbf{z}_{i,{T-P+1}:{T}};\mathbf{x}_{i,{T-P+1}:{T+Q}}\right),$ (2)
where $i=1,\dots,N$ denotes the identifier of a particular time series. Since the model outputs forecasts for the entire prediction horizon directly, it is also called a Seq2Seq model. Without loss of generality, we use the first approach as an example to illustrate our method, since both approaches focus on estimating a multivariate distribution $p\left(\mathbf{z}_{t}\right)$ or $p\left(\mathbf{z}_{i,{T+1}:{T+Q}}\right)$ (Fig. 1).
Figure 1: Illustration of the multivariate autoregressive and univariate Seq2Seq forecasting tasks.
A typical probabilistic forecasting model assumes Gaussian noise; for example, it models $\mathbf{z}_{t}$ as jointly following a multivariate Gaussian distribution:
$\left.\mathbf{z}_{t}\mid\mathbf{h}_{t}\right.\sim\mathcal{N}\left(\boldsymbol{\mu}(\mathbf{h}_{t}),\boldsymbol{\Sigma}(\mathbf{h}_{t})\right),$ (3)
where $\boldsymbol{\mu}(\cdot)$ and $\boldsymbol{\Sigma}(\cdot)$ are the functions mapping $\mathbf{h}_{t}$ to the mean and covariance parameters. The log-likelihood of the distribution given observed time series data up to time point $T$ can be used as the loss function for optimizing a DL model
$\mathcal{L}=\sum_{t=1}^{T}\log p\left(\mathbf{z}_{t}\mid\theta\left(\mathbf{h}_{t}\right)\right)\\ \propto\sum_{t=1}^{T}-\frac{1}{2}[\ln\lvert\boldsymbol{\Sigma}_{t}\rvert+\boldsymbol{\eta}_{t}^{\top}\boldsymbol{\Sigma}_{t}^{-1}\boldsymbol{\eta}_{t}],$ (4)
where $\boldsymbol{\eta}_{t}=\mathbf{z}_{t}-\boldsymbol{\mu}_{t}$. The above formulation simplifies to the univariate case when we set $N=1$ for the model, with the same model being shared across all time series:
$\left.z_{i,t}\mid\mathbf{h}_{i,t}\right.\sim\mathcal{N}\left(\mu(\mathbf{h}_{i,t}),\sigma^{2}(\mathbf{h}_{i,t})\right),$ (5)
where $\mu(\cdot)$ and $\sigma(\cdot)$ map $\mathbf{h}_{i,t}$ to the mean and standard deviation of a Gaussian distribution. The corresponding log-likelihood becomes
$\mathcal{L}=\sum_{t=1}^{T}\sum_{i=1}^{N}\log p\left(z_{i,t}\mid\theta\left(\mathbf{h}_{i,t}\right)\right)\propto\sum_{t=1}^{T}\sum_{i=1}^{N}-\frac{1}{2}\epsilon_{i,t}^{2}-\ln\sigma_{i,t},$ (6)
where $\epsilon_{i,t}=\frac{z_{i,t}-\mu_{i,t}}{\sigma_{i,t}}$. Eq. (4) and Eq. (6), when used as scoring rules to optimize the model, are generally referred to as the log-score and are widely employed in probabilistic forecasting.
For univariate problems, the CRPS is also a strictly proper scoring rule, defined as
$\operatorname{CRPS}(F,z)=\mathbb{E}_{F}|Z-z|-\frac{1}{2}\mathbb{E}_{F}\left|Z-Z^{\prime}\right|,$ (7)
where $F$ is the CDF of the predicted variable, $z$ is the observation, and $Z$ and $Z^{\prime}$ are independent random variables both associated with the CDF $F$. The CRPS has a closed-form expression when evaluating a Gaussian-distributed variable $z\sim\mathcal{N}\left(\mu,\sigma^{2}\right)$ (Gneiting et al.,, 2005):
$\operatorname{CRPS}\left(F,z\right)=z\left(2F\left(z\right)-1\right)+2f\left(z\right)-\frac{1}{\sqrt{\pi}},$ (8)
$\operatorname{CRPS}\left(F_{\mu,\sigma},z\right)=\sigma\operatorname{CRPS}\left(F,\frac{z-\mu}{\sigma}\right),$ (9)
where $F_{\mu,\sigma}\left(z\right)=F\left(\frac{z-\mu}{\sigma}\right)$, $F$ and $f$ are the CDF and PDF of the standard Gaussian distribution.
The CRPS has been shown to be a more robust alternative to the log-score as a loss function in univariate problems, particularly because the negative log-likelihood in Eq. (6) is quadratic in form (Gneiting et al.,, 2005; Rasp and Lerch,, 2018; Murad et al.,, 2021; Chen et al.,, 2024). Moreover, the CRPS can directly replace the log-score, providing analytical gradients with respect to $\mu$ and $\sigma$ for backpropagation. However, for a MVG distribution, the CRPS does not have a widely used closed-form expression.
3.2 MVG-CRPS as Loss Function for Multivariate Forecasting
In multivariate probabilistic forecasting, proper scoring rules such as the log-score (Eq. (4)) and the energy score ($\operatorname{ES}$) are used to evaluate predictive performance. The energy score generalizes CRPS to assess probabilistic forecasts of vector-valued random variables (Gneiting and Raftery,, 2007):
$\operatorname{ES}(F,\mathbf{z})=\mathop{\operatorname{\mathbb{E}}}_{\boldsymbol{Z}\sim F}\lVert\boldsymbol{Z}-\mathbf{z}\rVert^{\beta}-\frac{1}{2}\mathop{\operatorname{\mathbb{E}}}_{\begin{subarray}{c}\boldsymbol{Z}\sim F\\ \boldsymbol{Z}^{\prime}\sim F\end{subarray}}\lVert\boldsymbol{Z}-\boldsymbol{Z}^{\prime}\rVert^{\beta},$ (10)
[…]
$\operatorname{ES}(F,\mathbf{z})=\frac{1}{n}\sum_{i=1}^{n}\lVert\boldsymbol{Z}_{i}-\mathbf{z}\rVert^{\beta}-\frac{1}{2n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}\lVert\boldsymbol{Z}_{i}-\boldsymbol{Z}^{\prime}_{j}\rVert^{\beta},$ (11)
However, a significant disadvantage of using Eq. (11)) as the loss function is that it requires constant sampling during the training process, which can substantially slow down training.
In this section, we propose MVG-CRPS, a robust and efficient loss function designed as an alternative for multivariate forecasting. This loss function grows linearly with the prediction error, making it more robust than the log-score. Additionally, it does not require sampling during the training process, rendering it more efficient than the energy score.
Our proposed method is based on the whitening transformation of a time series vector that follows a MVG distribution, $\mathbf{z}_{t}\sim\mathcal{N}\left(\boldsymbol{\mu}_{t},\boldsymbol{\Sigma}_{t}\right)$. The whitening process transforms a random vector with a known covariance matrix into a new random vector whose covariance matrix is the identity matrix. As a result, the elements of the transformed vector have unit variance and are uncorrelated. This transformation begins by performing the singular value decomposition (SVD) of the covariance matrix:
$\boldsymbol{\Sigma}_{t}=\boldsymbol{U}_{t}\boldsymbol{S}_{t}\boldsymbol{U}_{t}^{\top},$ (12)
where $\boldsymbol{S}_{t}=\operatorname{diag}(\{\lambda_{t}^{1},\ldots,\lambda_{t}^{N}\})$ is a diagonal matrix containing the eigenvalues $\lambda_{t}^{i}$ of $\boldsymbol{\Sigma}_{t}$, and $\boldsymbol{U}_{t}$ is the orthonormal matrix of corresponding eigenvectors. We then define
$\mathbf{v}_{t}=\boldsymbol{U}_{t}^{\top}\left(\mathbf{z}_{t}-\boldsymbol{\mu}_{t}\right),$ (13)
where $\mathbf{v}_{t}\sim\mathcal{N}\left(\boldsymbol{0},\boldsymbol{S}_{t}\right)$ is a random vector with a decorrelated MVG distribution, having variances $\lambda_{i}$ (i.e., the corresponding eigenvalue) along the diagonal of its covariance matrix. Next, we define
$\mathbf{w}_{t}=\boldsymbol{S}_{t}^{-\frac{1}{2}}\mathbf{v}_{t}=\boldsymbol{S}_{t}^{-\frac{1}{2}}\boldsymbol{U}_{t}^{\top}\left(\mathbf{z}_{t}-\boldsymbol{\mu}_{t}\right),$ (14)
where $\mathbf{w}_{t}$ is a random vector with each element following a standard Gaussian distribution. We can then apply Eq. (8) individually to each element and obtain a new loss function mimicking Eq. (9) for multivariate problem:
$\mathcal{L}=\sum_{t=1}^{T}\sum_{i=1}^{N}\sqrt{\lambda_{t}^{i}}\operatorname{CRPS}\left(F,{w}_{i,t}\right).$ (15)
The primary advantage of this loss function lies in its ability to leverage the closed-form expression of $\operatorname{CRPS}$ by decorrelating the time series variables through whitening. While a potential limitation is that $\operatorname{CRPS}$, when applied to each transformed (whitened) variable, primarily evaluates the marginal distributions in the transformed space, the whitening process itself is derived from the original covariance matrix. As a result, the optimization process can still capture the covariance structure of the original distribution. Consequently, the loss function implicitly addresses certain aspects of the covariance, allowing $\operatorname{CRPS}$ to retain indirect sensitivity to the joint behavior of the original variables.
4 EXPERIMENTS
4.1 Datasets and Models
We apply the proposed MVG-CRPS to two multivariate forecasting tasks: multivariate autoregressive forecasting and univariate Seq2Seq forecasting. For the first task, we employ two benchmark models: the RNN-based GPVar (Salinas et al.,, 2019) and a decoder-only Transformer (Radford et al.,, 2018). For the second task, we use the N-HiTS model, which is based on multi-layer perceptrons (MLPs) (Challu et al.,, 2023).
To generate the distribution parameters for probabilistic forecasting, we employ a Gaussian distribution head based on the hidden state $\mathbf{h}_{i,t}$ produced by the model. Specifically, for the multivariate autoregressive forecasting, following Salinas et al., (2019), we parameterize the mean vector as $\boldsymbol{\mu}\left(\mathbf{h}_{t}\right)=\left[\mu_{1}\left(\mathbf{h}_{1,t}\right),\dots,\mu_{N}\left(\mathbf{h}_{N,t}\right)\right]^{\top}\in\mathbb{R}^{N}$ and adopt a low-rank-plus-diagonal parameterization of the covariance matrix $\boldsymbol{\Sigma}\left(\mathbf{h}_{t}\right)=\operatorname{diag}{(\mathbf{d}_{t})}+\boldsymbol{L}_{t}\boldsymbol{L}_{t}^{\top}$, where $\mathbf{d}_{t}=[d_{1}\left(\mathbf{h}_{1,t}\right),\ldots,d_{N}\left(\mathbf{h}_{N,t}\right)]^{\top}\in\mathbb{R}_{+}^{N}$ and $\boldsymbol{L}_{t}=\left[\mathbf{l}_{1}\left(\mathbf{h}_{1,t}\right),\dots,\mathbf{l}_{N}\left(\mathbf{h}_{N,t}\right)\right]^{\top}\in\mathbb{R}^{N\times R}$, $R\ll N$ is the rank parameter. Here, $\mu_{i}(\cdot)$, $d_{i}(\cdot)$, and $\mathbf{l}_{i}(\cdot)$ are the mapping functions that generate the mean and covariance parameters for each time series $i$ based on the hidden state $\mathbf{h}_{i={1:N},t}$. In practice, we use shared mapping functions across all time series, denoted as $\mu_{i}=\tilde{\mu}$, $d_{i}=\tilde{d}$, and $\mathbf{l}_{i}=\tilde{\mathbf{l}}$. This approach enables the use of random subsets of time series (i.e., batch size $B\leq N$) for model optimization in each iteration, making it feasible to apply our method to high-dimensional time series datasets. Similarly, in the univariate Seq2Seq forecasting task, the mean $\boldsymbol{\mu}\left(\mathbf{h}_{i}\right)$ and covariance $\boldsymbol{\Sigma}(\mathbf{h}_{i})$ are defined over the forecast horizon for each specific time series, based on the hidden states $\mathbf{h}_{i,t=T+1:T+Q}$. As a result, we can model the joint distribution $p\left(\mathbf{z}_{i,{T+1}:{T+Q}}\right)$ over the forecasted values.
We implemented our models using PyTorch Forecasting (Beitner,, 2020), with input data consisting of lagged time series values and covariates. We conducted extensive experiments on diverse real-world time series datasets obtained from GluonTS (Alexandrov et al.,, 2020). The prediction horizon ($Q$) and the number of rolling evaluations were taken from the configuration within GluonTS, where we follow the default setting by equating the context range to the prediction range, i.e., $P=Q$. Each dataset was sequentially split into training, validation, and testing sets. Each time series was individually normalized using a scaler fitted to its own training data. Predictions were then rescaled to their original values for computing evaluation metrics. For comprehensive details on the experimental setup, we refer readers to the Supplementary Materials (SM).
4.2 Quantitative Evaluation
We evaluate the proposed MVG-CRPS loss by comparing it with models trained using the log-score and energy score. The evaluation metrics are $\operatorname{CRPS}_{\text{sum}}$ and energy score (Eq. (11)). $\operatorname{CRPS}_{\text{sum}}$ is computed by summing both the forecast and ground-truth values across all time series and then calculating the $\operatorname{CRPS}$ (Eq. (7)) over the resulting sums (Salinas et al.,, 2019; Drouin et al.,, 2022; Ashok et al.,, 2024):
$\operatorname{\mathbb{E}}_{t}\left[\operatorname{CRPS}\left(F_{t},\sum_{i}z_{i,t}\right)\right],$ (16)
where the empirical $F_{t}$ is calculated by aggregating samples across time series, with 100 samples drawn to compute $\operatorname{CRPS}_{\text{sum}}$. The evaluation using the energy score is provided in the SM.
Table 1 presents a comparison of $\operatorname{CRPS}_{\text{sum}}$ across different scoring rules in the multivariate autoregressive forecasting task, underscoring the effectiveness of the proposed MVG-CRPS loss. MVG-CRPS consistently outperforms the log-score across most datasets. For instance, on the $\mathtt{electricity}$ and $\mathtt{tourism}$ datasets, MVG-CRPS achieves scores of 0.0249 and 0.2004, respectively, outperforming the scores of 0.0419 and 0.2217 achieved by the log-score. This indicates that MVG-CRPS leads to models that produce higher-quality forecasts. As discussed in later sections, this improvement is attributed to MVG-CRPS being less sensitive to outliers.
When compared to the energy score, MVG-CRPS demonstrates superior performance in 8 out of 13 datasets when using GPVar, and in 7 out of 13 datasets when using the Transformer model. For the remaining datasets, MVG-CRPS typically outperforms the log-score while maintaining comparable performance to the energy score, as highlighted in Table 1 (underlined). Additionally, as shown in Table 2, the energy score requires significantly more training time compared to both the log-score and MVG-CRPS. Although MVG-CRPS has a slightly longer overall training time than the log-score, it is faster per epoch, suggesting that it may require more epochs to converge but is computationally efficient per iteration.
The results for the univariate Seq2Seq forecasting task are presented in Table 3 and align with the findings from the multivariate autoregressive task. Overall, these findings demonstrate that MVG-CRPS provides a more accurate alternative to the log-score and a more efficient one compared to the energy score.
Table 1: Comparison of $\operatorname{CRPS}_{\text{sum}}$ across different scoring rules in the multivariate autoregressive forecasting task. The best scores are in boldface. Scores for MVG-CRPS are underlined when they are not the best overall but are better than the log-score. Mean and standard deviation are obtained from 10 runs of each model.
VAR GPVar Transformer
log-score energy score MVG-CRPS log-score energy score MVG-CRPS
$\mathtt{elec\_au}$ N/A 0.1261$\pm$0.0009 0.0887$\pm$0.0004 0.0967$\pm$0.0008 0.1633$\pm$0.0005 0.1492$\pm$0.0006 0.0793$\pm$0.0004
$\mathtt{cif\_2016}$ 1.0000$\pm$0.0000 0.0122$\pm$0.0004 0.0420$\pm$0.0006 0.0111$\pm$0.0005 0.0118$\pm$0.0003 0.0240$\pm$0.0014 0.0107$\pm$0.0002
$\mathtt{electricity}$ 0.1315$\pm$0.0006 0.0419$\pm$0.0008 0.0616$\pm$0.0004 0.0249$\pm$0.0006 0.0362$\pm$0.0002 0.0368$\pm$0.0004 0.0294$\pm$0.0004
$\mathtt{elec\_weekly}$ 0.1126$\pm$0.0011 0.1515$\pm$0.0028 0.0417$\pm$0.0014 0.0772$\pm$0.0031 0.0937$\pm$0.0026 0.0403$\pm$0.0013 0.0448$\pm$0.0014
$\mathtt{exchange\_rate}$ 0.0033$\pm$0.0000 0.0207$\pm$0.0004 0.0030$\pm$0.0001 0.0041$\pm$0.0001 0.0047$\pm$0.0003 0.0067$\pm$0.0003 0.0091$\pm$0.0004
$\mathtt{kdd\_cup}$ N/A 0.3743$\pm$0.0019 0.3210$\pm$0.0019 0.2358$\pm$0.0014 0.2076$\pm$0.0013 0.4789$\pm$0.0030 0.1959$\pm$0.0017
$\mathtt{m1\_yearly}$ N/A 0.4397$\pm$0.0041 0.4801$\pm$0.0022 0.3566$\pm$0.0029 0.5344$\pm$0.0109 0.3291$\pm$0.0047 0.4563$\pm$0.0111
$\mathtt{m3\_yearly}$ N/A 0.3607$\pm$0.0084 0.2186$\pm$0.0042 0.1423$\pm$0.0053 0.3156$\pm$0.0102 0.4050$\pm$0.0061 0.2325$\pm$0.0094
$\mathtt{nn5\_daily}$ 0.2303$\pm$0.0005 0.0998$\pm$0.0004 0.0958$\pm$0.0003 0.0948$\pm$0.0003 0.0991$\pm$0.0003 0.0883$\pm$0.0004 0.0811$\pm$0.0002
$\mathtt{saugeenday}$ N/A 0.4040$\pm$0.0047 0.3733$\pm$0.0048 0.3941$\pm$0.0055 0.3771$\pm$0.0088 0.3689$\pm$0.0053 0.3705$\pm$0.0047
$\mathtt{sunspot}$ N/A 18.7115$\pm$1.3296 23.3988$\pm$0.9662 17.2438$\pm$0.5833 39.7454$\pm$1.4841 16.6556$\pm$0.6167 22.6495$\pm$0.6752
$\mathtt{tourism}$ 0.1394$\pm$0.0012 0.2217$\pm$0.0027 0.2112$\pm$0.0014 0.2004$\pm$0.0022 0.2100$\pm$0.0017 0.2087$\pm$0.0020 0.2082$\pm$0.0015
$\mathtt{traffic}$ 3.5241$\pm$0.0084 0.0742$\pm$0.0004 0.0505$\pm$0.0002 0.0868$\pm$0.0002 0.0658$\pm$0.0002 0.0667$\pm$0.0002 0.0683$\pm$0.0000
Table 2: Comparison of training cost (in minutes) on GPVar across different scoring rules in the multivariate autoregressive forecasting task.
log-score energy score MVG-CRPS
per epoch total per epoch total per epoch total
$\mathtt{elec\_au}$ 0.86 33.53 16.29 717 0.78 29.14
$\mathtt{cif\_2016}$ 0.13 1.577 4.83 401.04 0.12 3.85
$\mathtt{electricity}$ 0.40 67.38 11.17 782.4 0.38 22.7
$\mathtt{elec\_weekly}$ 0.30 14.61 10.95 383.52 0.26 18.77
$\mathtt{exchange\_rate}$ 0.25 16.4 10.20 663.6 0.29 23.63
$\mathtt{kdd\_cup}$ 0.42 11.32 14.23 2063.52 0.42 28.79
$\mathtt{m1\_yearly}$ 0.19 3.708 5.66 469.92 0.18 8.019
$\mathtt{m3\_yearly}$ 0.43 7.299 10.80 291.72 0.42 14.49
$\mathtt{nn5\_daily}$ 0.29 9.208 11.64 244.5 0.27 14.53
$\mathtt{saugeenday}$ 0.23 12.65 10.70 524.46 0.15 15.32
$\mathtt{sunspot}$ 0.44 26.85 10.73 397.26 0.42 16.96
$\mathtt{tourism}$ 0.49 23.96 10.56 243 0.46 12.51
$\mathtt{traffic}$ 0.94 76.98 14.92 1044.6 0.92 92.46
Table 3: Comparison of $\operatorname{CRPS}_{\text{sum}}$ across different scoring rules in the univariate Seq2Seq forecasting task.
N-HiTS
log-score energy score MVG-CRPS
$\mathtt{covid}$ 0.1297$\pm$0.0048 N/A 0.1011$\pm$0.0022
$\mathtt{elec\_hourly}$ 0.0470$\pm$0.0008 N/A 0.0398$\pm$0.0004
$\mathtt{electricity}$ 0.0409$\pm$0.0003 0.0378$\pm$0.0006 0.0372$\pm$0.0003
$\mathtt{exchange\_rate}$ 0.0089$\pm$0.0005 0.0060$\pm$0.0002 0.0053$\pm$0.0002
$\mathtt{m4\_hourly}$ 0.0649$\pm$0.0007 0.0595$\pm$0.0005 0.0399$\pm$0.0007
$\mathtt{nn5\_daily}$ 0.0571$\pm$0.0003 0.0876$\pm$0.0006 0.0569$\pm$0.0004
$\mathtt{pedestrian}$ 0.7985$\pm$0.0511 0.9110$\pm$0.0210 0.5296$\pm$0.0071
$\mathtt{saugeenday}$ 0.4804$\pm$0.0150 0.4372$\pm$0.0100 0.3864$\pm$0.0035
$\mathtt{taxi\_30min}$ 0.0496$\pm$0.0002 0.0603$\pm$0.0002 0.0449$\pm$0.0001
$\mathtt{traffic}$ 0.2065$\pm$0.0007 0.0815$\pm$0.0001 0.0832$\pm$0.0002
$\mathtt{uber\_hourly}$ 0.7027$\pm$0.0209 0.6461$\pm$0.0052 0.5380$\pm$0.0033
$\mathtt{wiki}$ 0.0660$\pm$0.0011 0.0429$\pm$0.0003 0.0465$\pm$0.0004
4.3 Qualitative Evaluation
To further demonstrate the robustness and effectiveness of the MVG-CRPS loss, we compare the output covariance matrices from models trained with both the log-score and MVG-CRPS, and visualize the corresponding probabilistic forecasts.
4.3.1 Multivariate Autoregressive Forecasting
We first examine the multivariate autoregressive forecasting task using the $\mathtt{elec\_weekly}$ dataset (Fig. 2). When the model is trained with the log-score, the resulting covariance matrices exhibit disproportional variance values and occasionally display large covariance values. This observation is counterintuitive, given that the data have been normalized using scalers fitted to each time series. Such behavior can arise when large errors in the tails of the data disproportionately influence model training, causing the model to produce larger variances and covariances to accommodate these errors. In contrast, the model trained with MVG-CRPS produces covariance matrices with a more balanced distribution of variances and covariances. This indicates that MVG-CRPS effectively mitigates the influence of outliers, resulting in more reliable estimations of the predictive distribution.
To further illustrate the practical implications of our observations, we compare the probabilistic forecasts generated by GPVar on the $\mathtt{electricity}$ dataset (Fig. 3). The forecasts from the MVG-CRPS-trained model are noticeably more calibrated and sharper, especially when the predictions are projected to more reliable outcomes. In contrast, the model trained with the log-score occasionally produces much wider prediction intervals than MVG-CRPS, highlighting its susceptibility to outliers and resulting in less reliable forecasts (e.g., TS 1 in Fig. 3).
Figure 2: Comparison of the output covariance matrices $\boldsymbol{\Sigma}_{t}$ produced by GPVar on the $\mathtt{elec\_weekly}$ dataset. Top row and bottom row show covariance matrices from the model trained using the log-score and MVG-CRPS, respectively. For visual clarity, covariance values are clipped between 0 and 0.6.
Figure 3: Comparison of probabilistic forecasts produced by GPVar on the $\mathtt{electricity}$ dataset. The top row shows the forecasts from the model trained using the log-score, while the bottom row from the model trained using MVG-CRPS.
4.3.2 Univariate Seq2Seq Forecasting
The results for the univariate Seq2Seq forecasting task are presented in Fig. 4. Consistent with the findings from the multivariate autoregressive task, the model trained with the log-score tends to produce higher variance and covariance values. Consequently, the log-score-trained model may exhibit increased uncertainty, which can degrade the reliability of its forecasts. Fig. 4 illustrates the covariance over the prediction horizon of one day in the hourly $\mathtt{traffic}$ dataset. The covariance matrices directly model the prediction uncertainty over time, influenced by both the prediction lead time and the time of day. Uncertainty is generally larger during rush hours and increases with longer lead times. Compared to log-score-trained model, the MVG-CRPS-trained model is less affected by tail values while still effectively capturing these temporal patterns. This demonstrates that MVG-CRPS maintains the ability to model relevant uncertainty trends without being unduly influenced by outliers.
Another example is illustrated in Fig. 5, which showcases the probabilistic forecasts generated by the N-HiTS model on the $\mathtt{m4\_hourly}$ dataset. We observe notable differences in both the sharpness and calibration of the forecasts. Specifically, the model trained with MVG-CRPS produces slightly narrower prediction intervals, indicating greater confidence in its predictions compared to the log-score-trained model. In contrast, the log-score-trained model produces unnecessarily wide prediction intervals for simple time series with strong cyclical patterns. Additionally, the MVG-CRPS-trained model demonstrates higher accuracy in longer-term forecasts.
These qualitative evaluations underscore the advantages of using MVG-CRPS as a loss function for training probabilistic forecasting models. By reducing sensitivity to outliers and providing more stable covariance estimates, MVG-CRPS enhances the model to generate accurate and reliable probabilistic forecasts.
Figure 4: Comparison of the output covariance matrices $\boldsymbol{\Sigma}_{i}$ produced by N-HiTS on the $\mathtt{traffic}$ dataset. For visual clarity, covariance values are clipped between 0 and 1.0.
Figure 5: Comparison of probabilistic forecasts produced by N-HiTS on the $\mathtt{m4\_hourly}$ dataset.
5 CONCLUSION
In this paper, we have introduced MVG-CRPS, a robust loss function specifically designed for multivariate Gaussian outputs in probabilistic forecasting. Addressing the high sensitivity of the log-score to outliers and the inefficiency of the energy score, our approach leverages the CRPS to enhance robustness without compromising computational efficiency. The closed-form expression of MVG-CRPS based on neural network outputs facilitates its seamless integration into deep learning models. Our experimental results on real-world datasets demonstrate that MVG-CRPS outperforms existing loss functions like log-score and energy score, offering improved accuracy and robustness in probabilistic forecasting tasks. While we have focused on forecasting applications, the generality of our method allows it to be extended to a wide range of probabilistic regression problems involving multivariate Gaussian distributed responses.
For future work, we plan to explore the use of copulas to transform continuous distributions into Gaussian ones, thereby broadening the applicability of our approach to non-Gaussian settings. Additionally, we aim to investigate better parameterizations for covariance matrices to further enhance model performance and scalability. Overall, MVG-CRPS provides a general solution for multivariate Gaussian distribution forecasting, contributing to the development of more robust and accurate probabilistic models in the presence of outliers.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [No]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Not Applicable]
    2. (b)
      Complete proofs of all theoretical results. [Not Applicable]
    3. (c)
      Clear explanations of any assumptions. [Not Applicable]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes]
    2. (b)
      The license information of the assets, if applicable. [Not Applicable]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Not Applicable]
    4. (d)
      Information about consent from data providers/curators. [Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]